### v1

@@ 2.5 Code Stability
Figure 3: Comparison between our results and analogous simulations carried out by Leinhardt & Stewart (2012). Each panel shows results for collision outcome between a projectile and a target with mass ratio $m_{p}/m_{t}$ for a specific collision angle $\theta$. Larger $\theta$ values are for grazing collisions, while $\theta=0$ represents an head-on collision. The x-axis shows the ratio of the total kinetic over total potential energies in the center of mass reference frame.
In order for the code to remain stable and conserve energy, a sufficiently large elasticity constant and a sufficiently small time step need to be used. The requirements are that (i) the maximum compression $\Delta$ attainable in the simulation needs to be a small fraction $\epsilon_{\Delta}$ of the radius of the involved particles, and that (ii) at each time step $\delta{t}$ the movement of the fastest particle is only a small fraction $\epsilon_{t}$ of the maximum compression $\Delta$. For two particles $i$ and $j$ colliding with velocity $v_{i}$ and $v_{j}$ in their centers of mass, the compression can be written as:
$\Delta_{ij}=\frac{1}{2}\frac{m_{i}v_{i}^{2}}{k}\left(1+\frac{m_{i}}{m_{j}}\right)$ (10)
In order to ensure stability (i.e., that particles cannot penetrate each-other) we require:
@@
If we assume that $i$ is the lighter particle, then $\left(1+\frac{m_{i}}{m_{j}}\right)\leq 2$ and $\left(1+\frac{r_{j}}{r_{i}}\right)\geq 2$ and we conclude that the condition on the elasticity constant is:
@@
where $r_{\max}$ is the radius of the biggest particle in the simulation and $v_{\max}$ is the speed of the fastest particle in the simulation. In most cases, the above condition yields a minimum constant for physical reliability that is much smaller than the actual Young modulus of rocks, which is of the order of $10^{11}$ barye.
The condition on the time step is inextricably linked to the chosen elasticity constant, since larger constant yields more stiff particles and therefore smaller compression, which require a smaller time step to handle. The condition we enforce is
@@
where $m_{\min}$ is the mass of the smallest particle in the simulation.
It is to be noted that in simple binary collisions the actual value of the elasticity constant adopted affects predominantly the duration of a particle-particle bounce. It has, instead, a negligible impact on the forces involved and on the collision outcome. It is therefore not necessary to use the actual elasticity of the material in order to obtain reliable results. For that reason, it is more productive to select a value that satisfies Eq. 13 and then use Eq. 15 to derive the required time step.
@@ 2.6 Rubble-pile Formation
For the purposes of this paper, a rubble-pile is defined as a cluster of particles in a stable structure under self-gravity. To initialize such a cluster, particles are uniformly seeded in a spherical volume. A collision/overlap check is run across all particle pairs and if a collision/overlap is detected, one of the two particles is moved randomly to another location within the spherical domain limit. This process is repeated with a tolerance of 200 tries for all particles. If any collisions/overlaps are detected after 200 tries, the sphere radius is increased by the diameter of the largest particle. All particle positions are then randomly re-seeded in the new volume and the process is re-started. This eventually leads to a reasonably well packed cluster of particles (see the left panel of Figure 1). The loosely packed particles are then allowed to collapse under their own gravity into a stable structure. A rubble-pile formation is considered complete when the kinetic energy of the system has reduced to a negligible value compared to the collapsing phase. An example of a fully formed rubble pile is shown in the right panel of Figure 1.
@@ 2.7 Cluster-finding algorithm
A post-processing code was developed to compute the outcome of the collision in term of clusters of particles with long-term stability. This action is not trivial, since the simulation cannot be carried on within a reasonable time until all particles are either unbound or part of a stable cluster. The cluster finding algorithm works iteratively in three steps. A first step runs an adaptive Gaussian smoothing filter with a local standard deviation equal to ten times each particle radius and identifies the peak of the filtered mass distribution as the center of a potential stable cluster. The second step cycles through each particle that is not already part of the potential cluster deciding whether the particle is a member of the cluster by analyzing whether or not it is in contact with a particle that belongs to the cluster. This step is repeated until no additional contacts are identified. The third step cycles again through each particle that is not already part of the potential cluster deciding whether the particle is energetically bound to the cluster. The particle velocity with respect to the center of mass of the cluster is compared to the escape velocity at the position of the particle [^3] In this step we do not consider the direction of the relative velocity nor we investigate whether the particle is in collision course with another particle. Both these can potentially alter the belonging to the cluster. . If the latter is greater, the particle is accepted into the cluster. This step is also repeated until no new particle is accepted in the cluster. At this point the whole process is repeated on all the remaining particles that are unbound from all the previously found clusters. This cycle is repeated until no new cluster with more than a user-defined minimum number of particles is found.
@@
An example outcome is shown in Figure 2 where clusters and their member particles are color-coded. The case shown in the figure is the collision of two equal mass rubble piles with $\eta=3$. In this case six clusters with more than 50 particles are formed in the collision. The outer ones (1-green and 2-blue) are the biggest, with $M\sim 4.5\times 10^{22}$ g. The inner ones, instead, have smaller mass $M\sim 5\times 10^{21}$ g. The symmetry in the cluster properties is expected given the symmetry in the initial conditions. It has to be noted that very few of the cases in Figure 3 produce more than two cluster in the post-collision configuration. In most cases a single cluster is formed, except for the collisions between identical projectile-target, in which two similar clusters are most likely formed.
@@
Figure 4: Cartoon showing the collision geometry between a projectile (defined ad the object with lower mass) and a target (the object with higher mass). Collisions will be parameterized on their angle $\theta$. All collisions are carried out in the center of mass reference frame in which the velocities of the two objects are anti-parallel by construction.
3 Results
We used the DECCO code to carry out three experiments. We first run a set of validation collisions, in which we aim at replicating the initial conditions from Leinhardt & Stewart (2012), repeating their collision simulations, and eventually comparing our results with theirs. Once we verified the reliability of the code, we use DECCO’s capabilities to study the effect of rubble-pile collisions on the particle size distribution and the effect of pre-collision spin of the rubble piles on the collision outcome.
@@ 3.1 Validation Runs
We test the validity of our code by carrying out a set of simulated collisions between a pair of rubble-piles of different mass and colliding at different angles. These simulations are designed to emulate the analogous ones carried out by Leinhardt & Stewart (2012). We hold the target rubble pile mass (the largest of the pair) fixed, and carry out collisions with projectiles with mass one tenth of the target, one quarter of the target, and equal to the target. For each rubble pile pair, we carry out collisions with angles $\theta=0^{\circ}$, $22^{\circ}$, and $44^{\circ}$, as defined in Figure 4. For each rubble pile pair we also carry out collisions with different relative velocity, parameterized by the ratio of the kinetic energy involved over the binding potential energy of the two rubble piles. This parameter, which we call $\eta$, is defined as:
@@
All simulations were carried out with rubble piles with a particle distribution comprising three sizes: 7.5, 15, and 22.5 km. The total mass in each particle population was equal, resulting in most of the particles being of the smaller size. All particles had density $\rho=2.7$ g/cm3 (Flynn et al., 1999). The largest (target) rubble pile had 5000 total particles, while the two smaller projectiles had 1250 and 500 total particles, respectively. These choices resulted in rubble piles with mass and radius $(m_{\rm{RP}},r_{\rm{RP}})=(6\times 10^{21},95)$, $(10^{22},129)$, and $(6\times 10^{22},203)$, in units of g and km, respectively. Impact speed ranged between 0.26 km/s for the slowest equal-mass collision to $\sim 1$ km/s for the fastest collision with the light projectile. These impact velocities are smaller than the speed of sound in dense materials, which is of the order of a few km/s.
We adopted a time step $\delta{t}=0.04$ s and elastic constant $k_{\rm{in}}=10^{18}$, with a coefficient of restitution $0.9$. The elastic constant was chosen to respect Eq. 13 for the highest energy collisions and kept constant among different simulations (even if the maximum velocity was reduced) for consistency. The time interval also comfortably satisfies the constraint set by Eq. 15. The friction coefficients that we used were $\mu_{k}=0.3$ and $\mu_{r}=10^{-3}$.
The main difference between our calculations and the ones presented in Leinhardt & Stewart (2012) is the fact that we used rubble piles with a distribution of particle sizes, while Leinhardt & Stewart (2012) use mono-dispersed particles in their experiments. We compare the mass of the largest fragment in our simulation with that reported by Leinhardt & Stewart (2012) to evaluate the reliability of the code.
@@
While other authors have carried out similar calculations with different codes, we were not able to derive all the needed parameters and outcomes from their publications, and a direct comparison was not possible.
Based on the results of this comparison, we conclude that DECCO successfully compares with previously established results and is safe to use in the subsonic regime.
@@ 3.2 Particle size distribution
Having verified that the DECCO code’s prediction compare well with the result of other codes based on analogous physics and numerical techniques, we use the specific capability of DECCO to handle particles of different sizes to investigate the role of collisions in changing the particle distributions of rubble piles. In other words, we study whether the collisions predominantly eject the largest or the smallest particles in the system. This analysis is based on the same set of simulations that were used for the validation comparison of Fig. 3. The results of the analysis are shown in Figure 5. While there is significant scattering, there are also interesting patterns. First, any effect is small, at the level of a few per cent. Repeated collisions, however, could result on a build up making the effect more noticeable. Second, most constructive collisions (blue colored symbols) and low velocity collisions tend to preferentially eject small particles, leaving a debris characterized by a bigger particle population. This is especially true for asymmetric collisions (small projectile, squares), for which all constructive collisions result in a fragment with bigger average particle size. Most destructive collisions, instead, tend to eject bigger particles preferentially, with some notable exceptions at very high velocities.
@@
We further explore this in Figure 7. The figure is built to explore whether the high stresses are carried equally by particles of all sizes. The data shown are for an $\theta=22$ collision between a 1250 particle projectile and a 5000 particle target. Three values of the collision energy $\eta$ are shown. To create the figure, we first selected the top 1% of particles based on the strain that they are under (computed as the maximum particle overlap with all particles in contact). For each particle size $i$, the y-axis of Figure 7 shows the quantity $100\,N_{i,1\%}/N_{i}$, where $N_{i}$ is the number of particle of size $i$ in the entire simulation, and $N_{i,1\%}$ is the number of particles of size $i$ among the top $1\%$ strain particles. If all particle sizes were carrying strain equally, all y-axis values in the figures would be 1. That is not the case. As a general finding, the largest particles carry most of the strain at all times (the green line is always the largest). In addition, there is a clear evolution during and after the collision. In all cases, the pre-collision strain is emphasized with a horizontal dashed line. Let us first analyze the high-energy collision in the right panel. There is a short, few-minute spike during the collision (also shown in the right panel of Figure 6) in which large particles carry most of the strain, even more than in the pre-collision configuration. This is followed by a post-collision phase that lasts a few hours during which the rubble pile settles into a new configuration. In this phase, small grains increase their share of strain (green and orange lines are lower, blue line is higher). Eventually, the collision debris settle in a new configuration in which the largest particles still carry most of the strain but to a lesser extent with respect to the initial configuration. A similar pattern is observed for a medium energy collision in the central panel, albeit with reduced amplitude of the oscillations.
Figure 6: Volume rendering of the rubble piles before and during their collision. All particles are shown as a gray haze. Particles subject to the highest amount of stress are highlighted in dark gray. The collision is a constructive, low-velocity simulation ($\eta=3$) for a mass ratio 0.25 system with impact angle 22 degrees. The stresses are concentrated in the center of the target before the collision but the impact forces transfer the stresses to the area where the impact is happening, creating high stress chains that are carried predominantly by the bigger particles.
Figure 7: Strain analysis among the different particle sizes for a $\theta=22$, $m_{p}/m_{t}=0.25$ set of collisions of increasing energy (from left to right, $\eta=0.4$, $3.0$, and $12$). As detailed in the text, values larger than unity denote particle sizes that carry a larger than equal amount of the stress. The figure shows that large particles always carry the largest share of the stress to counter self-gravity in the clusters, and that this imbalance is even more pronounced during high-velocity impacts.

### v2

@@ 2.5 Code Stability
Figure 3: Comparison between our results and analogous simulations carried out by Leinhardt & Stewart (2012). Each panel shows results for collision outcome between a projectile and a target with mass ratio $m_{p}/m_{t}$ for a specific collision angle $\theta$. Larger $\theta$ values are for grazing collisions, while $\theta=0^{\circ}$ represents a head-on collision. The x-axis shows the ratio of the total kinetic over total potential energies in the center of mass reference frame. The y-axis, instead, shows the ratio between the mass of the largest fragment ($m_{\rm{lf}}$) and the mass of the target cluster ($m_{\rm{t}}$). See also Figure 5 for a definition of the collision angle $\theta$.
In order for the code to remain stable and conserve energy, a sufficiently large elastic constant and a sufficiently small time step need to be used. We follow Herrmann & Luding (1998) to derive constraints on the time step and elasticity constant. The requirements are that (i) the maximum compression $\Delta$ attainable in the simulation needs to be a small fraction $\epsilon_{\Delta}$ of the radius of the involved particles, and that (ii) at each time step $\delta{t}$ the movement of the fastest particle is only a small fraction $\epsilon_{t}$ of the maximum compression $\Delta$. For two particles $i$ and $j$ colliding with velocity $v_{i}$ and $v_{j}$ in their centers of mass, the compression can be written as:
$\Delta_{ij}=\sqrt{\frac{1}{2}\frac{m_{i}v_{i}^{2}}{k}\left(1+\frac{m_{i}}{m_{j}}\right)}$ (12)
In order to ensure stability (i.e., that particles cannot excessively penetrate each-other) we require:
@@
where $\rho$ is the particle density (see Table 1 for a list of parameters used). If we assume that $i$ is the lighter particle, then $\left(1+\frac{m_{i}}{m_{j}}\right)\leq 2$ and $\left(1+\frac{r_{j}}{r_{i}}\right)\geq 2$ and we conclude that the condition on the elasticity constant is:
@@
where $r_{\max}$ is the radius of the biggest particle in the simulation and $v_{\max}$ is the speed of the fastest particle in the simulation. In most cases, the above condition yields a minimum constant for physical reliability that is much smaller than the actual Young modulus of rocks, which is of the order of $10^{11}$ barye ($10^{10}$ N/m2).
The condition on the time step is inextricably linked to the chosen elasticity constant, since a larger constant yields more stiff particles and therefore smaller compressions, which require a smaller time step to handle. The condition we enforce is
@@
where $m_{\min}$ is the mass of the smallest particle in the simulation. We note that this is equivalent to imposing a time step that is much shorter than the oscillation period of a spring with an elastic constant $k$ loaded with the mass of the lightest particle considered.
It should be noted that in simple binary collisions, the actual value of the elasticity constant used predominantly affects the duration of a particle-particle bounce. It has a negligible impact on the forces involved and on the result of the collision [^4] The elastic constant, however, being directly related to the material’s Young modulus, does affect the speed of sound in the particles. Depending on the application, it may not be necessary to use the actual elasticity of the material in order to obtain reliable results. For that reason, it is more productive to select a value that satisfies Eq. 15 and then use Eq. 17 to derive the required time step. In our case, since we are interested in collisions at speed smaller than but close to the speed of sound, the adopted elasticity constant is similar to the physical one.
@@ 2.6 Rubble-pile Formation
In astronomy, a rubble pile is an object made of individual particles that have accumulated under the effect of self-gravity, for example after an energetic collision that has shattered a previously unfractured object (e.g., Bagatin et al. 2001). In addition, no single particle should contain more than half the total mass of the system. To initialize such a cluster, particles are uniformly seeded in a spherical volume. A collision/overlap check is run across all particle pairs and if a collision/overlap is detected, one of the two particles is moved randomly to another location within the spherical domain limit. This process is repeated with a tolerance of 200 tries for all particles. If any collisions/overlaps are detected after 200 tries, the sphere radius is increased by the diameter of the largest particle. All particle positions are then randomly re-seeded in the new volume and the process is re-started. This eventually leads to a reasonably well packed cluster of particles (see the left panel of Figure 1). The loosely packed particles are then allowed to collapse under their own gravity into a stable structure. Rubble-pile formation is considered complete when the kinetic energy of the system has reduced to a negligible value compared to the collapsing phase and to the collision energy involved in the subsequent simulation. All rubble piles used were relaxed until their kinetic energy was less than $10^{-6}$ the energy involved in the fastest collision, equivalent to less than $10^{-5}$ times their gravitational binding energy. An example of a fully formed rubble pile is shown in the right panel of Figure 1. Table 1 reports a list of symbols and, when relevant, the values that were adopted in our simulations.
We used the procedure described above to assemble asteroids with three sizes, the particle radii ranging from 95 to 203 km. The number of particles in the three was 500, 1250, and 5000. Each asteroid is made up of particles of three different sizes such that each particle size makes up a third of the overall mass of the asteroid. See also Sect 3 for further details.
@@ 2.7 Cluster-finding algorithm
A post-processing code was developed to compute the outcome of the collision in term of clusters of particles with long-term stability. This action is not trivial, since the simulation cannot be carried on within a reasonable time until all particles are either unbound or part of a stable cluster. The cluster finding algorithm works iteratively in three steps. A first step runs an adaptive Gaussian smoothing filter with a local standard deviation equal to ten times each particle radius and identifies the peak of the filtered mass distribution as the center of a potential stable cluster. The second step cycles through each particle that is not already part of the potential cluster deciding whether the particle is a member of the cluster by analyzing whether or not it is in contact with a particle that belongs to the cluster. This step is repeated until no additional contacts are identified. The third step cycles again through each particle that is not already part of the potential cluster deciding whether the particle is energetically bound to the cluster. The particle velocity with respect to the center of mass of the cluster is compared to the escape velocity at the position of the particle [^5] In this step we adopt a two-body energy criterion, and do not consider the direction of the relative velocity. We also do not investigate whether the particle is on a collision course with another particle. All of these can potentially alter the cluster to which a particle is bound to. . If the latter is greater, the particle is accepted into the cluster. This step is repeated until no new particles are accepted into the cluster. At this point the whole process is repeated on all the remaining particles that are unbound from all the previously found clusters. This cycle is repeated until no new cluster with more than a user-defined minimum number of particles is found.
Figure 4: Momentum conservation test for the DECCO code. The figure shows the result of a $\theta=22^{\circ}$ high-velocity collision between two large clusters, each with 5000 particles. The total kinetic energy in the collision is 12 times larger than the potential energy of the two clusters. The vertical axis shows the net momentum along each direction normalized by the total momentum carried by one of the clusters.
@@
An example outcome is shown in Figure 2 where clusters and their member particles are color-coded. The case shown in the figure is the collision of two equal mass rubble piles with three times more translational kinetic than potential energy ($\eta=3$, see below for a formal definition of the parameter $\eta$). In this case six clusters with more than 50 particles are formed in the collision. The outer ones (1-green and 2-blue) are the biggest, with $M\sim 4.5\times 10^{22}$ g. The inner ones, instead, have smaller mass $M\sim 5\times 10^{21}$ g. The symmetry in the cluster properties is expected given the symmetry in the initial conditions. It has to be noted that very few of the cases in Figure 3 produce more than two clusters in the post-collision configuration. In most cases a single cluster is formed, except for collisions between an identical projectile and target, in which two similar clusters are most likely formed.
@@
Figure 5: Cartoon showing the collision geometry between a projectile (defined as the object with lower mass) and a target (the object with higher mass). Collisions will be parameterized on their angle $\theta$. All collisions are carried out in the center of mass reference frame in which the velocities of the two objects are anti-parallel by construction.
3 Code testing and validation
An initial testing of the code was carried out to ensure that basic physics is preserved by testing for conservation of energy (in the case of unity coefficient of restitution) and momentum (for both unity and non-unity coefficient of restitution). All such tests were successful. We show in Figure 4 the result for momentum conservation in an oblique collision ($\theta=22^{\circ}$, as defined in Figure 5). The collision was carried out at the highest velocity (12 times more kinetic than potential energy) between two large clusters (5000 particles each). Figure 4 shows that momentum is conserved in all three spatial components to better than in one part per million.
In addition to conservation tests, we run a set of validation collisions, in which we aim at replicating the initial conditions from Leinhardt & Stewart (2012), repeating their collision simulations, and eventually comparing our results with theirs. We point out, however, that the code used by Leinhardt & Stewart (2012) is a hard-sphere code, while DECCO uses a soft-sphere method. In addition, Leinhardt & Stewart (2012) adopts a much smaller coefficient of restitution (0.5) compared to our value (0.9).
@@ 3.1 Validation Runs
We test the validity of our code by carrying out a set of simulated collisions between a pair of rubble-piles of different mass and colliding at different angles. These simulations are designed to emulate the analogous ones carried out by Leinhardt & Stewart (2012). We hold the target rubble pile mass (the largest of the pair) fixed, and carry out collisions with projectiles with mass one tenth of the target, one quarter of the target, and equal to the target. For each rubble pile pair, we carry out collisions with angles $\theta=0^{\circ}$, $22^{\circ}$, and $44^{\circ}$, as defined in Figure 5. For each rubble pile pair we also carry out collisions with different relative velocity, parameterized by the ratio of the translational kinetic energy involved over the binding potential energy of the two rubble piles. This parameter, which we call $\eta$, is defined as:
@@
All simulations were carried out with rubble piles with a particle distribution comprising three sizes: 7.5, 15, and 22.5 km. The total mass in each particle population was equal, resulting in most of the particles being of the smaller size. All particles had density $\rho=2.7$ g/cm3 (Flynn et al., 1999). The largest (target) rubble pile had 5000 total particles, while the two smaller projectiles had 1250 and 500 total particles, respectively. These choices resulted in rubble piles with mass and radius $(m_{\rm{RP}},r_{\rm{RP}})=(6\times 10^{21},95)$, $(10^{22},129)$, and $(6\times 10^{22},203)$, in units of g and km, respectively. Impact speed ranged between 0.26 km/s for the slowest equal-mass collision to $\sim 1$ km/s for the fastest collision with the light projectile. These impact velocities are smaller than the speed of sound in dense materials, which is of the order of a few km/s. However, they may exceed the speed of a wave’s propagation through a granular medium at low confining pressure (Sánchez et al., 2022). In addition, some particles might break upon impact, a capability that is not yet implemented in DECCO. All our collision velocities $v_{\rm{coll}}$ exceeded the system escape velocity $v_{\rm{esc}}$, ranging from $v_{\rm{coll}}\simeq\,14v_{\rm{esc}}$ in the slowest case to $v_{\rm{coll}}\simeq 250\,v_{\rm{esc}}$.
We adopted a time step $\delta{t}=0.04$ s and elastic constant $k_{\rm{in}}=10^{18}$, with a coefficient of restitution [^6] This is a somewhat high value adapted from the measurement of Durda et al. (2011). They found a coefficient of restitution of 0.83 that we round up conservatively to 0.9. $\epsilon=0.9$. The elastic constant was chosen to respect Eq. 15 for the highest energy collisions and kept constant among different simulations (even if the maximum velocity was reduced) for consistency. The time interval also comfortably satisfies the constraint set by Eq. 17. The friction coefficients that we used were $\mu_{k}=0.3$ and $\mu_{r}=10^{-3}$.
As indicated above, there are some significant differences between our calculations and the ones presented in Leinhardt & Stewart (2012). First, the version of PKDGRAV used in Leinhardt & Stewart (2012) is based on collisions between hard spheres, not soft spheres like in DECCO. Second, the coefficient of restitution that they adopt is significantly lower (0.5) than the value that we use. Finally, we used rubble piles with a distribution of particle sizes, while Leinhardt & Stewart (2012) use mono-dispersed particles in their experiments. Taking these caveats into account, we compare the mass of the largest fragment in our simulation with that reported by Leinhardt & Stewart (2012) to assess the reliability of DECCO.
@@
While other authors have carried out similar calculations with different codes, we were not able to derive all the needed parameters and outcomes from their publications, and a direct comparison was not possible. Based on the results of this comparison, we conclude that DECCO successfully compares with previously established results and is safe to use in the subsonic regime.
4 Results
@@ 4.1 Particle size distribution
Having verified that the DECCO code respects basic conservation laws and compares well with the result of other codes based on analogous physics, we use the specific capability of DECCO to handle particles of different sizes to investigate the role of collisions in changing the particle distributions of rubble piles. In other words, we study whether the collisions predominantly eject the largest or the smallest particles in the system. This analysis is based on the same set of simulations that were used for the validation comparison of Fig. 3. The results of the analysis are shown in Figure 6. While there is significant scattering, there are also interesting patterns. First, any effect is small, at the level of a few per cent. Repeated collisions, however, could result on a build up making the effect more noticeable. Second, most constructive collisions (blue colored symbols) and low velocity collisions tend to preferentially eject small particles, leaving a debris characterized by a bigger particle population. This is especially true for asymmetric collisions (small projectile, squares), for which all constructive collisions result in a fragment with bigger average particle size. Most destructive collisions, instead, tend to eject bigger particles preferentially, with some notable exceptions at very high velocities.
@@
We further explore this in Figure 8. The figure is built to explore whether the high stresses are carried equally by particles of all sizes. The data shown are for a $\theta=22^{\circ}$ collision between a 1250 particle projectile and a 5000 particle target. Three values of the collision energy $\eta$ are shown. To create the figure, we first selected the top 1% [^7] Note that the choice of the top 1% of particles is arbitrary. However, the behavior described in this section is robust against a different choice of what defines the high-stress category of particles. of particles based on the stress that they are under (computed as the maximum particle overlap with all particles in contact). For each particle size $i$, the y-axis of Figure 8 shows the quantity $100\,N_{i,1\%}/N_{i}$, where $N_{i}$ is the number of particle of size $i$ in the entire simulation, and $N_{i,1\%}$ is the number of particles of size $i$ with the top $1\%$ stress. If all particle sizes were equally under stress, all y-axis values in the figures would be 1. That is not the case. As a general finding, the largest particles carry most of the stress at all times (the green line is always the largest). In addition, there is a clear evolution during and after the collision. In all cases, the pre-collision stress is emphasized with a horizontal dashed line. Let us first analyze the high-energy collision in the right panel. There is a short, few-minute spike during the collision (also shown in the right panel of Figure 7) in which large particles carry most of the stress, even more than in the pre-collision configuration. This is followed by a post-collision phase (shown in Figure 8 for $5^{\prime}\lesssim{t}\lesssim 100^{\prime}$) that lasts a few hours during which the rubble pile settles into a new configuration. In this phase, small grains increase their share of stress (green and orange lines are lower, blue line is higher). Eventually, the collision debris settle in a new configuration in which the largest particles still carry most of the stress but to a lesser extent with respect to the initial configuration. A similar pattern is observed for a medium energy collision in the central panel, albeit with reduced amplitude of the oscillations.
Figure 7: Volume rendering of the rubble piles before and during their collision. All particles are shown as a gray haze. Particles subject to the highest amount of stress are highlighted in dark gray. The collision is a constructive, low-velocity simulation ($\eta=3$) for a mass ratio 0.25 system with an impact angle of 22 degrees. The stresses are concentrated in the center of the target before the collision but the impact forces transfer the stresses to the area where the impact is happening, creating high stress chains that are carried predominantly by the bigger particles.